If x = 2 is one solution of x^2 + 8x - c = 0, then the other solution is…
(A)20 (B) 10 (C) - 2 (D) - 10 (E) - 20

Answer by stanbon(75887)

If x = 2 is one solution of x^2 + 8x - c = 0, then the other solution is…
(A)20 (B) 10 (C) - 2 (D) - 10 (E) - 20
------------------
Use synthetic division:
2)....1....8....-c
.......1....10...|..-c+20
Solve -c+20 = 0
c = 20
-----
Now solve x^2+8x-20 = 0
Factor:
(x+10)(x-2)=0
x = -10 or x = 2
==================
Cheers,
Stan H.